OpenStudy (anonymous):

What is the simplified form of x minus 2 over x squared plus x minus 6 divided by x squared plus 5x plus 4 over x plus 4 ?

OpenStudy (ikram002p):

Grrr...

OpenStudy (ikram002p):

like this ? \(\Large x-\frac{2}{x^2}+{x-\frac{6}{x^2}}+\frac{4}{x}+4\)